OpenStudy (shaleiah):

@abtster @mathmath333 @TheSmartOne

OpenStudy (mathmath333):

no clue here

OpenStudy (mathmath333):

@surjithayer

OpenStudy (mathmath333):

@freckles

OpenStudy (freckles):

Can you type it here? It is blurry to me.

OpenStudy (freckles):

Well I can tell whatever thing is except that exponent.

OpenStudy (mathmath333):

yes it seems 2 or 3

OpenStudy (mathmath333):

its 2 cuz its quadratic stated at the top

OpenStudy (freckles):

y=x^3 has range all real numbers y=x^2 has range y>=0

OpenStudy (freckles):

alright

OpenStudy (freckles):

Well if you have y=-x^2 the range is y<=0

OpenStudy (freckles):

if you have y=-x^2+4 the range is y<=4

OpenStudy (freckles):

so if you have y=-ax^2+4 the range is y<=4 where a>0 constant

OpenStudy (freckles):

So look at your function...

OpenStudy (freckles):

And try to use what I have said here... If you want you can graph it.

OpenStudy (freckles):

so graph y=-ax^2+4 where a>0 we have